OpenStudy (anonymous):

For each matrix, determine if the matrix is diagonalizable, nilpotent or neither. If the matrix is diagonalizable, find a matrix P that diagonalizes the matrix. If the matrix is nilpotent, find a string basis associated with the matrix. If he matrix is neither, find the Jordan canonical form of the matrix. A = [5 4 2] [4 5 2] [2 2 2]

OpenStudy (mertsj):

I do not know the meaning of those terms.

OpenStudy (anonymous):

I can't find the eigenvalues!

OpenStudy (anonymous):

@Calcmathlete

OpenStudy (anonymous):

@apoorvk

OpenStudy (zarkon):

what is your characteristic polynomial?

OpenStudy (anonymous):

What do you mean?

OpenStudy (apoorvk):

Do you have any idea about what the Jordan canonical form is? I am familiar with other terms but not this one...

OpenStudy (anonymous):

No but we can find other terms

OpenStudy (zarkon):

the eigenvalues are easy to find...I'm not sure what your issue is.

OpenStudy (anonymous):

I don't know I got this equation: x^3 - 17x^2 + 25x - 26 = 0 I don't how to solve from here

OpenStudy (zarkon):

that is not what I get

OpenStudy (anonymous):

what do you get?

OpenStudy (zarkon):

\[-\lambda^3+12\lambda^2-21\lambda+10\]

OpenStudy (zarkon):

the above equation is the characteristic polynomial

OpenStudy (zarkon):

equal to zero that is

OpenStudy (anonymous):

I don't know how did you get this equation and how to solve it?

OpenStudy (zarkon):

I computed \[\det\left(A-\lambda I_3\right)\]

OpenStudy (anonymous):

I did the samething

OpenStudy (zarkon):

then you must have made a mistake in your calculation

OpenStudy (anonymous):

ok how to solve this equation?

OpenStudy (zarkon):

easy...it factors ...you get 1,1,10 as the roots

OpenStudy (anonymous):

it's cubic equation!

OpenStudy (zarkon):

so?

OpenStudy (anonymous):

how to solve it?!

OpenStudy (zarkon):

factor it...use the rational root test

OpenStudy (anonymous):

can you show me please?

OpenStudy (zarkon):

test \(\pm10,\pm5,\pm2,\pm1\) into the above equation you will see that 10 and 1 are the only ones that work...and in fact 1 is a repeated root.

OpenStudy (anonymous):

ok how did you get these numbers?

OpenStudy (anonymous):

but you said it's 1, 1 and 10

OpenStudy (zarkon):

yes..that is correct

OpenStudy (zarkon):

that is what I said both time I gave you the roots

OpenStudy (anonymous):

ok sorry

OpenStudy (anonymous):

So it's gonna be 1 and 10?

OpenStudy (zarkon):

yes...where 1 is a repeated root

OpenStudy (anonymous):

ok So it's not diagonalizable right?

OpenStudy (zarkon):

it is ;)

OpenStudy (anonymous):

how come ? it's not equal to n

OpenStudy (zarkon):

this is one of those weird cases where you have a repeated root but it is still diagonalize

OpenStudy (anonymous):

so it's considered 3 not 2?

OpenStudy (zarkon):

have you tried to find the eigenvectors?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Why do I want to find the eigenvector?

OpenStudy (zarkon):

have you ever diagonalized a matrix before?

OpenStudy (anonymous):

ok sorry got it! to find P right?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok I got 4b1 + 4b2 + 2b3 = 0

OpenStudy (anonymous):

for lambda = 1

OpenStudy (zarkon):

I get \[p=\left[\begin{matrix}-1 & 2 & -1 \\ 1 & 2 & 0\\ 0 & 1 & 2\end{matrix}\right]\]

OpenStudy (anonymous):

Can you explain how to get this?

OpenStudy (zarkon):

ok...can you solve your equation in terms of the free variables

OpenStudy (anonymous):

do you mean do it in rref?

OpenStudy (zarkon):

I assume you looked at A-I and put it in rref

OpenStudy (anonymous):

I don't how to get P

OpenStudy (zarkon):

that is what we are working on...there are several steps to go through

OpenStudy (zarkon):

Put \(A-I_3\) into rref I get \[\left[\begin{matrix}1 & 1 & 1/2 \\ 0 & 0 & 0\\ 0 & 0 & 0\end{matrix}\right]\]

OpenStudy (anonymous):

A-1 = [ 1 , 1 , 0.5 ; 0 , 0 , 0 ; 0 , 0 , 0]

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

now can you solve this system in terms of the free variables

OpenStudy (anonymous):

to get what?

OpenStudy (zarkon):

the eigenspace (the eigenvectors)

OpenStudy (anonymous):

i don't know?

OpenStudy (zarkon):

have you ever found a Null space before?

OpenStudy (anonymous):

yes, I got b1 = -b2 and b3 = -2 b1

OpenStudy (zarkon):

if we use x,y,z then [1 , 1 , 0.5] turns into x+y+.5z=0 let z=t and y=r they are free variables then we get the solution x=-r-.5t,y=y, z=t so we have \[\left(\begin{matrix}-r-.5t \\ r\\t\end{matrix}\right)=\left(\begin{matrix}-.5t \\ 0\\t\end{matrix}\right)+\left(\begin{matrix}-r \\ r\\0\end{matrix}\right)\] \[=\left(\begin{matrix}-.5 \\ 0\\1\end{matrix}\right)t+\left(\begin{matrix}-1 \\ 1\\0\end{matrix}\right)r\]

OpenStudy (zarkon):

typo *then we get the solution x=-r-.5t,y=r, z=t

OpenStudy (anonymous):

ok got it. This is the eignvector? right?

OpenStudy (zarkon):

\[\left(\begin{matrix}-.5 \\ 0\\1\end{matrix}\right)\text{ and }\left(\begin{matrix}-1 \\ 1\\0\end{matrix}\right)\] are though I used \[\left(\begin{matrix}-1 \\ 0\\2\end{matrix}\right)\text{ and }\left(\begin{matrix}-1 \\ 1\\0\end{matrix}\right)\]

OpenStudy (zarkon):

notice these two vectors are in the matrix P that I wrote above

OpenStudy (anonymous):

from where you got 2 2 1 ?

OpenStudy (zarkon):

that is the eigenvector corresponding to the eigenvalue 10

OpenStudy (anonymous):

are we working on 1 or 10 ?

OpenStudy (zarkon):

we just did 1...since it was a repeated root we got 2 vectors

OpenStudy (anonymous):

I don't get this part, so this is the eignvector for lambda = 1 right?

OpenStudy (zarkon):

there are two vectors for lambda =1. the two vectors I wrote above

OpenStudy (anonymous):

ok where is the third one comes from?

OpenStudy (zarkon):

from lambda =10

OpenStudy (anonymous):

so for lambda = 1 we have 2 vectors and for lambda = 10 we have 3 vectors?

OpenStudy (zarkon):

no...we get one vector from lambda =10. which gives us 3 total

OpenStudy (anonymous):

ohh ok got it Can you tell how to get the nilpoten?

OpenStudy (anonymous):

multiply the matrix to get the zero vector?

OpenStudy (zarkon):

it is not nilpotent

OpenStudy (anonymous):

So I don't have to find it?

OpenStudy (anonymous):

becasue the det(A) is not = 0 right?

OpenStudy (zarkon):

if A was nilpotent then \(A^n=0\) for some \(n\) what would imply that the determinant of A was zero. But the determinant of A is not zero. Thus A is not nilpotient

OpenStudy (zarkon):

correct

OpenStudy (anonymous):

ok so for this matrix the answer is done?

OpenStudy (anonymous):

because I have two more matrices and I will try to do it by myself

OpenStudy (zarkon):

you have what you need

OpenStudy (anonymous):

ok Thank you so much for your help and your time I really appreciate that

OpenStudy (zarkon):

No problem ... I just hope you undrstand some of it now

OpenStudy (anonymous):

yea I do